# A GREAT Architecture for Edge-Based Graph Problems Like TSP

### v2

@@ Abstract
In the last years, many learning-based approaches have been proposed to tackle combinatorial optimization problems such as routing problems. Many of these approaches are based on graph neural networks (GNNs) or related transformers, operating on the Euclidean coordinates representing the routing problems. However, models operating on Euclidean coordinates are ill-suited for non-Euclidean, asymmetric problem instances that are often found in real-world settings. To overcome this limitation, we propose a novel GNN-based and edge-focused neural model called Graph Edge Attention Network (GREAT). Using GREAT as an encoder to capture the properties of a routing problem instance, we build a reinforcement learning framework which we apply to Euclidean and non-Euclidean variants of vehicle routing problems such as Traveling Salesman Problem, Capacitated Vehicle Routing Problem and Orienteering Problem. Our framework is among the first to tackle non-Euclidean variants of these problems and achieves competitive results among learning-based solvers.
1 Introduction
Graph neural networks (GNNs) have emerged as a powerful tool for learning on graph-structured data such as molecules, social networks, or citation graphs (Wu et al. 2020). In recent years, GNNs have also been applied in the setting of combinatorial optimization, especially routing problems (Joshi, Laurent, and Bresson 2019; Hudson et al. 2021; Xin et al. 2021) since such problems can be modeled as graph problems. Alternatively, transformer models (Vaswani et al. 2017) have been used as encoder architectures for routing problems in learning-based settings (Kool, van Hoof, and Welling 2019; Kwon et al. 2020; Jin et al. 2023b). Both architectures, GNNs and transformers, are used to iteratively compute node embeddings of the problem instances. They do so through message-passing operations and the attention mechanism, respectively, starting from the Euclidean coordinates and, optionally, other node features like customer demands as initial features. However, as a result, these architectures are limited to Euclidean problem instances. In reality, routing problems are often non-Euclidean or even asymmetric due to, e.g., traffic congestion (and time being the objective), one way streets (and distance being the objective) or elevation changes (and energy consumption being the objective). Non-Euclidean routing problems are specified by pair-wise distances between the nodes (e.g., in form of distance matrices). Such pair-wise distances are examples of graph edge features and can not be represented as node features like Euclidean coordinates. As a result, since GNNs and transformers focus on operating on and computing node-level features, they are not well-suited for such asymmetric problems.
@@ 1 Introduction
  • We utilize GREAT in a reinforcement learning (RL) framework that can be trained end-to-end to predict optimal tours of the Traveling Salesman Problem (TSP), Capacitated Vehicle Routing Problem (CVRP) and Orienteering Problem (OP). As the inputs of GREAT are edge features (e.g., distances), GREAT applies to all variants of such routing problems, including non-Euclidean variants like the asymmetric TSP. The resulting trained framework achieves state-of-the-art performance for two different asymmetric distribution in TSP, CVRP and OP with 100 nodes. Notably, our approach outperforms MatNet (Kwon et al. 2021) which is the current standard for non-Euclidean problems.
  • We propose a few-shot curriculum learning (CL) framework to fine-tune our trained GREAT models from TSP with 100 nodes to bigger instance sizes. It is based on incrementally increasing the problem instance size up to 500 nodes and test the generalization performance of the obtained models. In an experimental evaluation, we show that GREAT has strong generalization performance (even in a zero-shot setting without CL) and achieves competitive performance for asymmetric TSP up to $1,000$ nodes.
2 Preliminaries
2.1 Graph Neural Networks
@@
GNNs iteratively compute node feature vectors by aggregating over the node feature vectors of adjacent nodes and mapping the old feature vector together with the aggregation to a new node feature vector. Additionally, the feature vectors are multiplied with trainable weight matrices and nonlinearities are applied to achieve actual learning. The node feature vectors of a neighborhood are typically scaled in some way (depending on the respective GNN architecture) and, sometimes, edge feature vectors are also considered within the aggregations. A conceptual example can be found in Figure 1. For the mathematical formulation of such an iterative node feature update, we consider the Graph Attention Network (GATs (Velickovic et al. 2017)) whose update formula is given by:
@@
While the original GAT does not incorporate edge features in its message passing operations, some extensions of the architecture integrate both node and edge features for improved representation learning. For example, Chen and Chen (2021) propose Edge-Featured Graph Attention Networks (EGAT) which uses edge features by applying a GAT not only on the input graph itself but also its line graph representation (compare Chen, Li, and Bruna (2017) as well) and then combining the computed features. Another work incorporating both node and edge features in the node-level aggregations in an attention-based GNN is Shi et al. (2020) who use a “Graph Transformer” for a semi-supervised classification task. Further, Jin et al. (2023a) introduce “EdgeFormers”, an architecture operating on Textual-Edge Networks where they combine the success of Transformers in LLM tasks and GNNs. Their architecture also augments GNNs to utilize edge (text) features. By this, text representations (e.g., the text of a product review corresponding to edge $(i,u)$ of an item $i$ by a user $u$) can be enriched by structural graph data. We note, however, that none of these architectures are purely edge-feature-focused.
@@
2.2 Learning to Route
@@
Popular approaches for solving routing problems with the help of machine learning include RL frameworks, where encodings for the problem instances are computed. These encodings are then used to incrementally build solutions by selecting one node in the problem instance at a time. Successful works in this category include Deudon et al. (2018); Nazari et al. (2018); Kool, van Hoof, and Welling (2019); Kwon et al. (2020); Jin et al. (2023b).
@@
Another possibility to use machine learning for solving routing problems is to predict edge probabilities or scores which are later used in search algorithms such as beam search or guided local search. Examples for such works are Joshi, Laurent, and Bresson (2019); Kool, van Hoof, and Welling (2019); Fu, Qiu, and Zha (2021); Xin et al. (2021); Hudson et al. (2021); Sun and Yang (2023); Min, Bai, and Gomes (2024).
@@
A further possibility involves iterative methods where a solution to a routing problem is improved over and over until a stopping criterion (e.g., convergence) is met. Possibilities for such improvements are optimizing subproblems or applying improvement operators such as $k$-opt. Examples for such works are da Costa et al. (2021); Wu et al. (2021); Cheng et al. (2023); Lu, Zhang, and Yang (2019); Chen and Tian (2019); Li, Yan, and Wu (2021).
Non-Euclidean Routing Problems
@@
Many of the mentioned works, especially in the first two categories, use GNNs or transformer models (which are related to GNNs via GATs (Joshi 2020)) to capture the structure of the routing problem in their neural architecture. This is done by interpreting the coordinates of Euclidean routing problem instances as node features. These node features are then processed in the GNN or transformer architectures to produce encodings of the overall problem. However, this limits the applicability of such works to Euclidean routing problems. This is unfortunate, as non-Euclidean routing problems are highly relevant in reality (Boyacı, Dang, and Letchford 2021). Consider, e.g., one-way streets which result in unequal travel distances between two points depending on the direction one is going. Another example is variants of TSP that consider energy consumption as the objective to be minimized. If point A is located at a higher altitude than point B, traveling from A to B might require less energy than the other way around. Furthermore, e.g. during rush-hours, travel times can be unequal between two points due to one-way directional congestion.
@@
So far, only a few studies have also investigated non-Euclidean versions of routing problems, such as the asymmetric TSP (ATSP). One such study is Gaile et al. (2022) where they solve synthetic ATSP instances with unsupervised learning, RL, and supervised learning approaches. Another study is Wang et al. (2023) that uses online RL to solve ATSP instances of TSPLIB (Reinelt 1991). Another successful work tackling ATSP is Kwon et al. (2021). There, the Matrix Encoding Network (MatNet) is proposed, a neural model suitable to operate on matrix encodings representing combinatorial optimization problems such as the distance matrices of (A)TSP. Their model is trained using RL. Ye et al. (2024) proposes a modular framework to solve larger routing problem instances and uses MatNet as a subsolver to tackle bigger (A)TSP instances. Drakulic et al. (2023) proposes a learning formulation that allows training on relatively small routing problem instances while still performing well on larger instances at inference. While doing so, they also tackle ATSP, using a “normal” GNN as their encoder that uses random node features as the inputs together with edge features that encode the actual distances.
3 Graph Edge Attention Network
@@
Figure 2: GREAT (node-free): edge attends to adjacent edges
We propose an edge-level-based GNN where information is passed along neighboring edges. This makes our model tailored for edge-level tasks such as edge classification (e.g., in the context of routing problems, determining if edges are “promising” to be part of the optimal solution or not). Our model is attention-based, meaning the “focus” of an edge to another adjacent edge in the update operation is determined using the attention mechanism. Consequently, similar to the Graph Attention Network (GAT) we call our architecture Graph Edge Attention Network (GREAT). A simple visualization of the idea of GREAT is shown in Figure 2. In this visualization, edge $e_{14}$ attends to all other edges it shares an endpoint with. While GREAT is a task-independent framework, it is suited perfectly for routing problems: Consider (A)TSP as an example. In this setting, node features are typically absent, only edge features are given as distances between nodes. A standard node-level GNN would not be well-suited to process such data. Existing work uses coordinates of the nodes in the Euclidean space as node features to overcome this limitation (e.g., (Joshi, Laurent, and Bresson 2019; Kool, van Hoof, and Welling 2019)). However, this strategy is limited to Euclidean TSP and does not extend to other forms of the problem, like asymmetric cases. GREAT, however, can be applied to all TSP variants.
While in theory it would also be possible to construct a line graph to transform edges into nodes and by this make “normal” GNNs applicable to the problem, this would transform an original dense (A)TSP graph with $n$ cities and $n^{2}$ edges into a representation with $\mathcal{O}(n^{2})$ nodes and $\mathcal{O}(n^{3})$ edges (Hudson et al. 2021), increasing the order of magnitude of the representation by one.
@@ 3.1 Architecture
The first variant is purely edge-focused and does not have any node features. Here, each edge exchanges information with all other edges it shares at least one endpoint with. The idea essentially corresponds to the visualization in Figure 2. In the following, we refer to this variant as “node-free” GREAT or GREAT “NF”.
The second variant is also edge-focused but has intermediate, temporary node features. This essentially means that nodes are used to save all information on adjacent edges. Afterwards, the features of an edge are computed by combining the temporary node features of their respective endpoints. These node features are then deleted and not passed on to the next layer, only the edge features are passed on. The idea of this GREAT variant is visualized in Figures 3 and 4. In the remainder of this study, we refer to this GREAT version as “node-based” or GREAT “NB”.
@@
The advantages and disadvantages of both GREAT variants focus mostly on technical implementation aspects. While both GREAT variants were implemented in Python using Pytorch (Paszke et al. 2019), the GREAT NB implementation requires additional message-passing operations by Pytorch Geometric (Fey and Lenssen 2019) which makes the architecture slightly slower. However, the GREAT NB implementation generalizes more easily to non-complete, sparse graphs. We note that the code (as well as the trained models and datasets of the experiments in Section 4) will be publicly available after the paper is accepted.
3.2 Mathematical Formulations
We now describe the mathematical equations defining the internal operations of GREAT. We note that, inspired by the original transformer architecture of (Vaswani et al. 2017), GREAT consists of two types of sublayers: attention sublayers and feedforward sublayers. We always alternate between attention and feedforward sublayers. The attention sublayers can be node-based (with temporary nodes features) or completely node-free. Using the respective sublayers leads to overall node-based or node-free GREAT. A visualization of the block stacking can be found in the Supplementary Material (SM).

### v3

@@ Abstract
In the last years, an increasing number of learning-based approaches have been proposed to tackle combinatorial optimization problems such as routing problems. Many of these approaches are based on graph neural networks (GNNs) or related transformers, operating on the Euclidean coordinates representing the routing problems. However, such models are ill-suited for a wide range of real-world problems that feature non-Euclidean and asymmetric edge costs. To overcome this limitation, we propose a novel GNN-based and edge-focused neural model called Graph Edge Attention Network (GREAT). Using GREAT as an encoder to capture the properties of a routing problem instance, we build a reinforcement learning framework which we apply to both Euclidean and non-Euclidean variants of vehicle routing problems such as Traveling Salesman Problem, Capacitated Vehicle Routing Problem and Orienteering Problem. Our framework is among the first to tackle non-Euclidean variants of these problems and achieves competitive results among learning-based benchmarks.
I Introduction
Graph neural networks (GNNs) have emerged as a powerful tool for learning on graph-structured data such as molecules, social networks, or citation graphs [1]. In recent years, GNNs have also been applied in the setting of combinatorial optimization, especially routing problems [2, 3, 4] since such problems can be modeled as graph problems. Alternatively, transformer models [5] have been used as encoder architectures for routing problems in learning-based settings [6, 7, 8]. Both architectures, GNNs and transformers, are used to iteratively compute node embeddings of the problem instances. They do so through message-passing operations and the attention mechanism, respectively, starting from the Euclidean coordinates and, optionally, other node features like customer demands as initial features. However, as a result, these architectures are limited to Euclidean problem instances.
In reality, routing problems are often non-Euclidean or even asymmetric due to, e.g., traffic congestion (and time being the objective), one way streets (and distance being the objective) or elevation changes (and energy consumption being the objective). Non-Euclidean routing problems are specified by pair-wise distances between the nodes (e.g., in form of distance matrices). Such pair-wise distances are examples of graph edge features and can not be represented as node features like Euclidean coordinates. As a result, since GNNs and transformers focus on operating on and computing node-level features, they are not well-suited for such asymmetric problems.
@@ I Introduction
  • We utilize GREAT in a reinforcement learning (RL) framework that can be trained end-to-end to predict optimal tours of the Traveling Salesman Problem (TSP), Capacitated Vehicle Routing Problem (CVRP) and Orienteering Problem (OP). As the inputs of GREAT are edge features (e.g., distances), GREAT applies to all variants of such routing problems, including non-Euclidean variants like the asymmetric TSP. The resulting trained framework achieves state-of-the-art performance for two different asymmetric distribution in TSP, CVRP and OP with 100 nodes. Notably, our approach outperforms MatNet [9] which is the current standard for non-Euclidean problems.
  • We propose a few-shot curriculum learning (CL) framework to fine-tune our trained GREAT models from TSP with 100 nodes to bigger instance sizes. It is based on incrementally increasing the problem instance size up to 500 nodes and, by this, familiarizing the obtained models with larger TSP instances.
  • In an additional experimental evaluation, we investigate the generalization performance of GREAT. We show that GREAT generalizes well to larger TSP instances (even in a zero-shot setting without CL) and achieves competitive performance for asymmetric TSP up to $1,000$ nodes. Further, we show that GREAT is also applicable to real-world routing problems, achieving strong generalization performance on TSPLIB, CVRPLIB and OPLIB.
The remainder of this article is structured as follows: we introduce relevant background and literature in Section II. Afterwards, GREAT is introduced in Section III. Our GREAT-based RL framework, the exact experimental setup and its results are presented in Section IV before we conclude in Section V.
II Preliminaries
II-A Graph Neural Networks
@@
GNNs iteratively compute node feature vectors by aggregating over the node feature vectors of adjacent nodes and mapping the old feature vector together with the aggregation to a new node feature vector. Additionally, the feature vectors are multiplied with trainable weight matrices and nonlinearities are applied to achieve actual learning. The node feature vectors of a neighborhood are typically scaled in some way (depending on the respective GNN architecture) and, sometimes, edge feature vectors are also considered within the aggregations. A conceptual example can be found in Figure 1. For the mathematical formulation of such an iterative node feature update, we consider the Graph Attention Network (GATs [10]) whose update formula is given by:
@@
While the original GAT does not incorporate edge features in its message passing operations, some extensions of the architecture integrate both node and edge features for improved representation learning. For example, [11] propose Edge-Featured Graph Attention Networks (EGAT) which uses edge features by applying a GAT not only on the input graph itself but also its line graph representation (compare [12] as well) and then combining the computed features. Another work incorporating both node and edge features in the node-level aggregations in an attention-based GNN is [13] who use a “Graph Transformer” for a semi-supervised classification task. Further, [14] introduce “EdgeFormers”, an architecture operating on Textual-Edge Networks where they combine the success of Transformers in LLM tasks and GNNs. Their architecture also augments GNNs to utilize edge (text) features. By this, text representations (e.g., the text of a product review corresponding to edge $(i,u)$ of an item $i$ by a user $u$) can be enriched by structural graph data. We note, however, that none of these architectures are purely edge-feature-focused.
@@
II-B Routing Problems
II-B1 Traveling Salesman Problem
The traveling salesman problem (TSP) is an NP-hard combinatorial optimization problem. Given a set of nodes (representing “cities”) and the pairwise distances between them, he objective of the problem is to find a shortest route that visits all nodes exactly once and returns to the starting node in the end. We provide an example of an Euclidean TSP (where the nodes are distributed in the Euclidean plane) including a possible solution tour in Figure 2(a). The depicted tour was generated using the well-known LKH heuristic [15]. In practice, TSP instances are often asymmetric. Therefore, in this paper, we consider not only Euclidean TSP instances (EUC; with coordinates sampled uniformly at random in the unit square) but additionally TSP instances of the TMAT distribution (following [9]). TMAT leads to asymmetric distances that follow the triangle inequality, which means that for three nodes $u,v,w$ in our set of nodes it holds that $d(u,v)\leq d(u,w)+d(w,v)$. We follow the TMAT distribution generation of [9] but normalize the distance matrices such that the biggest distance between two nodes is exactly 1. Further, we consider the XASY distribution (extremely asymmetric) where all pairwise distances are sampled uniformly at random from the range $(0,1)$ which means that the triangle inequality between different nodes does generally not hold, making the problem more challenging. The same distribution has been used before in [16].
II-B2 Capacitated Vehicle Routing Problem
In the capacitated vehicle routing problem (CVRP), we are given a set of customer nodes with associated demands. Further, we have a special depot node and a vehicle with a certain, fixed capacity. The goal of the problem is to find a route that starts and ends in the depot, visits all customers exactly once to deliver “goods” in accordance to their demands and minimizes the traveled distance. Since in general the total amount of customer demands exceeds the vehicle capacity, it is usually necessary to intermediately return to the depot several times. An example of a Euclidean CVRP instance with a possible tour can be found in Figure 2(b). The tour was generated using the well-known HGS heuristic [17]. In our work, we consider CVRP instances where the pairwise distances between nodes are sampled from the same EUC, TMAT and XASY distributions used for TSP. Further, we use the customer demands and vehicle capacity used in [6], which means for 100 customers we have a vehicle with a capacity of 50 and the demands are sampled uniformly at random from $\{1,\dots,9\}$.
II-B3 Orienteering Problem
In the orienteering problem (OP), each node (representing a customer) has a collectible prize. Further, we have a vehicle that starts in a designated depot location. The goal of the problem is to find a tour that maximizes the prices collected from the the customers while not traveling further than a specified maximum distance (and, again, it is not allowed to visit a customer more than once). In the end, the vehicle returns to the depot node. We provide an example of a Euclidean OP instance with 50 customers in Figure 2(c). The solution to the problem was found using the EA4OP heuristic [18]. In this work, the distances between the nodes in our OP instances follow the earlier introduced EUC, TMAT and XASY distributions. The prizes that can be collected by visiting the customers are sampled using a similar uniform specification to [6], i.e., we assign prizes sampled uniformly from $\{\frac{1}{100},\dots,\frac{100}{100}\}$ to our customers. The maximum travel distance (considering instances with 100 customers) is set to $4$ for EUC and TMAT instances and to $0.4$ for XASY instances.
(a) TSP instance with 50 nodes
(b) CVRP instance with 50 customers
(c) OP instance with 50 customers
Figure 2: Routing problem examples
II-C Learning to Route
@@
Popular approaches for solving routing problems with the help of machine learning include RL frameworks, where encodings for the problem instances are computed. These encodings are then used to incrementally build solutions by selecting one node in the problem instance at a time. Successful works in this category include [19, 20, 6, 7, 8].
@@
Another possibility to use machine learning for solving routing problems is to predict edge probabilities or scores which are later used in search algorithms such as beam search or guided local search. Examples for such works are [2, 6, 21, 4, 3, 22, 23].
@@
A further possibility involves iterative methods where a solution to a routing problem is improved over and over until a stopping criterion (e.g., convergence) is met. Possibilities for such improvements are optimizing subproblems or applying improvement operators such as $k$-opt. Examples for such works are [24, 25, 26, 27, 28, 29].
II-C1 Non-Euclidean Routing Problems
@@
Many of the mentioned works, especially in the first two categories, use GNNs or transformer models (which are related to GNNs via GATs [30]) to capture the structure of the routing problem in their neural architecture. This is done by interpreting the coordinates of Euclidean routing problem instances as node features. These node features are then processed in the GNN or transformer architectures to produce encodings of the overall problem. However, this limits the applicability of such works to Euclidean routing problems. This is unfortunate, as non-Euclidean routing problems are highly relevant in reality [31].
@@
So far, only a few studies have also investigated non-Euclidean versions of routing problems, such as the asymmetric TSP (ATSP). One such study is [16] where they solve synthetic ATSP instances with unsupervised learning, RL, and supervised learning approaches. Another study is [32] that uses online RL to solve ATSP instances of TSPLIB [33]. Another successful work tackling ATSP is [9]. There, the Matrix Encoding Network (MatNet) is proposed, a neural model suitable to operate on matrix encodings representing combinatorial optimization problems such as the distance matrices of (A)TSP. Their model is trained using RL. [34] proposes a modular framework to solve larger routing problem instances and uses MatNet as a subsolver to tackle bigger (A)TSP instances. [35, 36] proposes a learning formulation that allows training on relatively small routing problem instances while still performing well on larger instances at inference. Further, they propose a model that is trained for several routing problems at once. While doing so, they also tackle ATSP, using a “normal” GNN as their encoder that uses random one-hot vectors as node-feature inputs together with edge features that encode the actual distances.
III Graph Edge Attention Network
@@
Figure 3: GREAT (node-free): edge “attends” to adjacent edges
Existing GNNs are based on node-level message-passing operations. In contrast, we propose an edge-level-based GNN where information is compared between and passed along neighboring edges. This makes our model tailored for edge-level tasks such as edge classification (e.g., in the context of routing problems, determining if edges are “promising” to be part of the optimal solution or not). Our model is attention-related, meaning the “focus” of an edge to another adjacent edge in the update operation is determined using an attention-inspired mechanism. Consequently, similar to the Graph Attention Network (GAT) we call our architecture Graph Edge Attention Network (GREAT). A simple visualization of the idea of GREAT is shown in Figure 3. In this visualization, edge $e_{14}$ compares information between (or “attends” to) all other edges it shares an endpoint with. While GREAT is a task-independent framework, it is suited perfectly for routing problems: Consider (A)TSP as an example. In this setting, node features are typically absent, only edge features are given as distances between nodes. A standard node-level GNN would not be well-suited to process such data. Existing work uses coordinates of the nodes in the Euclidean space as node features to overcome this limitation (e.g., [2, 6]). However, this strategy is limited to Euclidean TSP and does not extend to other forms of the problem, like asymmetric cases. GREAT, however, can be applied to all TSP variants.
While in theory it would also be possible to construct a line graph to transform edges into nodes and by this make “normal” GNNs applicable to the problem, this would transform an original dense (A)TSP graph with $n$ cities and $n^{2}$ edges into a representation with $\mathcal{O}(n^{2})$ nodes and $\mathcal{O}(n^{3})$ edges (compare [3]), increasing the order of magnitude of the representation by one.
@@ III-A Architecture
The first variant is purely edge-focused and does not have any node features. Here, each edge compares information with all other edges it shares at least one endpoint with. The idea essentially corresponds to the visualization in Figure 3. In the following, we refer to this variant as “node-free” GREAT or GREAT “NF”.
The second variant is also edge-focused but has intermediate, temporary node features. This essentially means that nodes are used to save all information on adjacent edges. Afterwards, the features of an edge are computed by combining the temporary node features of their respective endpoints. These node features are then deleted and not passed on to the next layer, only the edge features are passed on. The idea of this GREAT variant is visualized in Figures 4 and 5. In the remainder of this study, we refer to this GREAT version as “node-based” or GREAT “NB”.
@@
The advantages and disadvantages of both GREAT variants focus mostly on technical implementation aspects. While both GREAT variants were implemented in Python using Pytorch [37], the GREAT NB implementation requires additional message-passing operations by Pytorch Geometric [38] which makes the architecture slightly slower. However, the GREAT NB implementation generalizes more easily to non-complete, sparse graphs. Moreover, we hypothesize that the NB variant ensures a more considerable information exchange between adjacent edges compared to the purely edge-based variant. We note that the code (as well as the trained models and datasets of the experiments in Section IV) are publicly available. [^1] https://github.com/attila-lischka/GREAT
III-B Mathematical Formulations
We now describe the mathematical equations defining the internal operations of GREAT. Similar to the original transformer architecture of [5], GREAT consists of two types of sublayers: attention sublayers and feedforward sublayers. We always alternate between attention and feedforward sublayers. The attention sublayers can be node-based (with temporary nodes features) or completely node-free. Using the respective sublayers leads to overall node-based or node-free GREAT. A visualization of the block stacking can be found in Figure 6.